OpenStudy (anonymous):

h(x)=square root ((1/x^2+7)+16 if f(x)=square root x, then find a function g such that h=f o g and, if f(x)=square root(x+16), then find a function g such that h=f o g.

OpenStudy (anonymous):

g(x)=(1/x^2+7)+16 for f(x)=sqrt(x) g(x)=(1/(x^2+7) for f(x)=sqrt(x+16)

OpenStudy (anonymous):

thank you